OpenStudy (moongazer):

(27)^x-1=(1/9)^x . Find x. Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

(27)^x-1=(1/9)^x can be rewritten as powers of 3 Hence (3^3) ^x-1 = (3^-2) ^x Multiply powers Hence 3^3(x-1) = 3^-2x Equate powers 3(x-1) = -2x 3x - 3 = -2x add 2x to both sides 5x - 3 = 0 add 3 to both sides 5x = 3 divide both sides by 5 x = 3/5 Any misunderstanding just ask

OpenStudy (moongazer):

Thanks! i understood it. :)

OpenStudy (moongazer):

could answer my other question?